# The shocks in Josephson transmission line revisited

### 27 Dec 2023 v3

@@ D.1 Integration by quadrature
Consider the particular case $p_{3}=p_{1}$. In this case (D) can be easily integrated by quadrature, provided the coefficient before the first derivative is equal to zero, i.e. $m$ is
$m=\sqrt{\frac{\alpha}{n+1}}p_{1}.$ (67)
Note that the condition of compatibility of (67) and (64) (the latter with $p_{3}=p_{1}$) is
$\gamma=\sqrt{\frac{\alpha}{n+1}}(n+2)p_{1}.$ (68)
@@
$w_{z}^{2}-\frac{1}{n^{2}p_{1}^{2}}w^{2n+2}=2E,$ (70)
where $E$ is the integration constant. So finally, the equation
$p_{\tau\tau}+\gamma p_{\tau}=\alpha p\left(p^{2n}-p_{1}^{2}\right)$ (71)
can be integrated by quadrature when the coefficients of the equation satisfy Eq. (68). Note that we could have decreased the number of independent parameters in (71) by introducing the new dependent variable $\overline{p}=p/p_{1}^{1/n}$, after which the equation would take the form
$\overline{p}_{\tau\tau}+\gamma\overline{p}_{\tau}=\alpha p_{1}^{2}\overline{p}\left(\overline{p}^{2n}-1\right).$ (72)
@@ D.2 The elementary solutions
$\lim_{\tau\to-\infty}p(\tau)=p_{1}^{1/n},\hskip 28.45274pt\lim_{\tau\to+\infty}p(\tau)=0.$ (73)
Hence the solution of (70) should satisfy the boundary condition
$\lim_{z\to\infty}z^{1/n}w(z)=p_{1},$ (74)
@@
$w(z)=\frac{p_{1}^{1/n}}{\left(1+z\right)^{1/n}}.$ (75)
@@
$p(\tau)=\frac{p_{1}^{1/n}.}{\left[\exp\left(\sqrt{\frac{\alpha}{n+1}}np_{1}\tau\right)+1\right]^{1/n}},$ (76)
which exists provided the constants of (71) are connected by the relation (68).
Now we have a pleasant surprise: the elementary solution (75) solves (D) also when $p_{3}\neq p_{1}$. In fact, substituting (75) into (D) we obtain
$\frac{(n+1)m^{2}}{(1+z)^{2}}-\frac{(n+1)m^{2}-\alpha p_{1}p_{3}}{z(1+z)}$
$=\frac{\alpha p_{1}^{2}}{(1+z)^{2}}-\frac{\alpha(p_{1}-p_{3})p_{1}}{z(1+z)},$ (77)
and mysteriously the terms in the l.h.s and in the r.h.s. of the equation are equal pairwise, provided $m$ is given by (67). Hence (76) is the elementary solution of (61) valid also when $p_{3}\neq p_{1}$). The condition of compatibility of (64) and (67) in the general case is
$\gamma=\sqrt{\frac{\alpha}{n+1}}[p_{1}+(n+1)p_{3}]$ (78)
@@
A few remarks concerning the obtained solution would be we believe relevant. First, because of the symmetry of (61) we could have obtained Eqs. (76) and (78) with $p_{1}$ and $p_{3}$ interchanged and $p$ changed to $-p$. The choice between the two particular solutions was made by imposing the boundary conditions (73) (the first one). Second, if we want that $p(\tau)$ exists and remains finite for $\tau\in(-\infty,+\infty)$ and want to impose the boundary conditions at $\tau=\pm\infty$, Eq. (73) is the only possible choice (apart from changing $p_{1}$ to $p_{3}$). The trajectory should start in the unstable fixed point and end in the stable fixed point.
D.2.1 The strongly asymmetric case
For $p_{3}\gg p_{1}$ there exists additional elementary solution of (61). In this case the equation can be approximated as
$p_{\tau\tau}+\gamma p_{\tau}=\alpha p_{3}p\left(p^{n}-p_{1}\right).$ (79)
Comparing (79) with (71) (and keeping in mind the elementary solution of the latter (76)) we obtain the elementary solution of (79) in the form
$p(\tau)=\frac{p_{1}^{1/n}}{\left\{\exp\left[\sqrt{\frac{\alpha p_{3}p_{1}}{2(n+2)}}n\tau\right]+1\right\}^{2/n}}.$ (80)
where $\gamma$ and $\alpha$ are connected by the equation (see (68))
$\gamma=\sqrt{\frac{\alpha p_{3}p_{1}}{2(n+2)}}(n+4).$ (81)
Looking at Eq. (61) we realise that for $p_{3}\gg p_{1}$, the condition (78) corresponds to strong damping, when the second derivative term in the l.h.s. of the equation is small in comparison with the first derivative term, and the condition (81) – to critical damping, when both terms are of the same order of magnitude.
D.2.2 Abel equation
The obtained elementary solutions become more transparent after we notice that introducing new dependent variable $y\equiv p_{\tau}$ and considering $p$ as the independent variable, we may write down (61) as Abel equation of the second kind
$yy_{p}+\gamma y=\alpha p\left(p^{n}-p_{1}\right)\left(p^{n}+p_{3}\right).$ (82)
For $\gamma$ and $\alpha$ connected by the formula (78), Eq. (82) obviously has the particular solution
$p_{\tau}=\sqrt{\frac{\alpha}{n+1}}p\left(p^{n}-p_{1}\right)$ (83)
(we substituted $p_{\tau}$ for $y$ in the l.h.s. of the equation). Integrating thus obtained differential equation we obtain (76).
Approximating (82) for $p_{3}\gg p_{1}$ as
$yy_{p}+\gamma y=\alpha p_{3}p\left(p^{n}-p_{1}\right)$ (84)
we notice that for $\gamma$ and $\alpha$ connected by the formula (81), Eq. (84) has the particular solution
$p_{\tau}=\sqrt{\frac{2\alpha p_{3}}{n+2}}p\left(p^{n/2}-p_{1}^{1/2}\right)$ (85)
(we substituted $p_{\tau}$ for $y$ in the l.h.s. of the equation). Integrating thus obtained differential equation we obtain (80).
It is curious that Eq. (82) doesn’t help much in obtaining the general solution of (61) which, on the other hand, can easily be obtained by integrating (70). Actually, it is the other way round – this general solution can be used to obtain parametric representation of the general solution of Abel equation (under the conditions $p_{3}=p_{1}$ and (68) polyanin .

### 28 Feb 2024 v6

@@ D.1 Integration by quadrature
Consider the particular case $x_{3}=x_{1}$. In this case (D) can be easily integrated by quadrature, provided the coefficient before the first derivative is equal to zero, i.e. $m$ is
$m=\sqrt{\frac{\alpha}{n+1}}x_{1}.$ (76)
Note that the condition of compatibility of (76) and (73) (the latter with $x_{3}=x_{1}$) is
$\gamma=(n+2)m.$ (77)
@@
$w_{z}^{2}-\frac{1}{n^{2}x_{1}^{2}}w^{2n+2}=2E,$ (79)
where $E$ is the integration constant.
@@ D.2 The elementary solutions
$\lim_{\tau\to-\infty}x(\tau)=x_{1}^{1/n},\hskip 28.45274pt\lim_{\tau\to+\infty}x(\tau)=0.$ (80)
Note that if the solution of (70) $x(\tau)$ should exist and remain finite for $\tau\in(-\infty,+\infty)$, Eq. (80) are the only possible boundary conditions at $\tau=\pm\infty$ (apart from changing $x_{1}$ to $x_{3}$). The trajectory should start in the infinite past in the unstable fixed point and end in the infinite future in the stable fixed point.
From (80) follows that the solution of (79) should satisfy the boundary condition
$\lim_{z\to\infty}z^{1/n}w(z)=x_{1},$ (81)
@@
$w(z)=\frac{x_{1}^{1/n}}{\left(1+z\right)^{1/n}}.$ (82)
@@
$x(\tau)=\frac{x_{1}^{1/n}.}{\left[\exp\left(nm\tau\right)+1\right]^{1/n}},$ (83)
which exists, provided the constants of (70) are connected by the relation (77).
Now we have a pleasant surprise: the elementary solution (82) solves (D) also when $x_{3}\neq x_{1}$. In fact, substituting (82) into (D) we obtain
$\frac{(n+1)m^{2}}{(1+z)^{2}}-\frac{(n+1)m^{2}-\alpha x_{1}x_{3}}{z(1+z)}$
$=\frac{\alpha x_{1}^{2}}{(1+z)^{2}}-\frac{\alpha(x_{1}-x_{3})x_{1}}{z(1+z)},$ (84)
and mysteriously the terms in the l.h.s and in the r.h.s. of the equation are equal pairwise, provided $m$ is given by (76). Hence (83) is the elementary solution of (70) valid also when $x_{3}\neq x_{1}$). The condition of compatibility of (73) and (76) in the general case is
$\gamma=\sqrt{\frac{\alpha}{n+1}}[x_{1}+(n+1)x_{3}]$ (85)
@@
D.3 The strongly asymmetric case
For $x_{3}\gg x_{1}$ there exists additional elementary solution of (70). In this case the equation can be approximated as
$x_{\tau\tau}+\gamma x_{\tau}=\alpha x_{3}x\left(x^{n}-x_{1}\right)$
$=\alpha x_{3}x\left(x^{n/2}-x_{1}^{1/2}\right)\left(x^{n/2}+x_{1}^{1/2}\right).$ (86)
Comparing (D.3) with (70) (and keeping in mind the elementary solution of the latter (83), we obtain the elementary solution of (D.3)
$x(\tau)=\frac{x_{1}^{1/n}}{\left[\exp\left(nm^{\prime}\tau\right)+1\right]^{2/n}}.$ (87)
where
[…]
D.4 Abel equation
The obtained elementary solutions become more transparent after we notice that introducing new dependent variable $p$
$p=x_{\tau}$ (89)
and considering $x$ as the independent variable, we may write down (70) as Abel equation of the second kind polyanin
$pp_{x}+\gamma p=\alpha x\left(x^{n}-x_{1}\right)\left(x^{n}+x_{3}\right).$ (90)
For $\gamma$ and $\alpha$ connected by the formula (85), Eq. (90) has the particular solution
$p=\sqrt{\frac{\alpha}{n+1}}x\left(x^{n}-x_{1}\right)$ (91)
Substituting $p(x)$ into (89) and integrating thus obtained differential equation we obtain (83).
Approximating (90) for $x_{3}\gg x_{1}$ as
$pp_{x}+\gamma p$ $=$ $\alpha x_{3}x\left(x^{n}-x_{1}\right)$
$=$ $\alpha x_{3}x\left(x^{n/2}-x_{1}^{1/2}\right)\left(x^{n/2}+x_{1}^{1/2}\right),$
we obtain from (91) the particular solution of (D.4)
$p=\sqrt{\frac{2\alpha x_{3}}{n+2}}x\left(x^{n/2}-x_{1}^{1/2}\right),$ (93)
and from (85) the connection between $\gamma$ and $\alpha$ in the present case; the latter turns out to be identical to that given by (88). Substituting obtained $p(x)$ into (89) and integrating thus obtained differential equation we obtain (87).